### v1

Abstract
The hypothesis of the relation between the observed alignment of spots in the X-ray films in cosmic ray emulsion experiments and the selection procedure of the highest-energy particles itself together with the transverse momentum conservation is tested in the framework of the HYDJET++ model. It is shown that the high degree of alignment can appear at the reasonable values of transverse momentum disbalance of selected most energetic particles.
1 Introduction
In our previous study Lokhtin:2023tze ; LokhtinYF24 in the framework of geometrical approach it has been shown that the high degree of alignment can appear partly due to the selection procedure of the highest-energy particles itself and the threshold on the energy deposition together with the transverse momentum conservation. This hopeful result encourages us to investigate more intently this hypothesis in the framework of some realistic model of hadron interactions. It is about the observation of strong angular correlations in the families of hadrons and photons that are products of interaction between cosmic rays and target nuclei pamir1 ; pamir2 ; pamir3 ; pamir4 ; book . This phenomenon can potentially indicate the coplanar structure of the observed events and was named alignment pamir4 . At the qualitative level, alignment shows a deviation from the straight line of most energetic registered particles and their clusters in the plane of emulsion film. In addition to that the strong angular correlations of particles was also observed in stratospheric strat1 ; strat2 X-ray-emulsion chamber experiments. So far there is no commonly accepted explanation of this intriguing phenomenon in spite of numerous attempts to find it (see, for instance, book ; halzen ; man ; mukh ; Lokhtin:2005bb ; Lokhtin:2006xv ; deroeck and references therein).
[…]
2 Basic definitions and kinematical relations
To characterize the azimuthal angular correlations the Pamir Collaboration defines the alignment of spots in term of the variable introduced by A. Borisov pamir4 :
$\lambda_{N}~{}=~{}\frac{\sum^{N}_{i\neq j\neq k}\cos(2\phi_{ijk})}{N(N-1)(N-2)},$ (1)
where $\phi_{ijk}$ is the angle between two straight lines connecting the $i$spot with the the $j$th and $k$th spots. $N$ is the number of spots used in an analysis. This parameter $\lambda_{N}$ describes the degree of alignment better than the possible other parameters of asymmetry like the eccentricity or the thrust. For instance, $\lambda_{4}$ will be equal to 1 if all four points lie strictly on the same straight line, but it will be considerably less than 1 if these points form four vertices of a long rectangle. The combinatorial factor $(N(N-1)(N-2))$ normalizes the parameter $\lambda_{N}$, because it takes into account the number of variants of choosing three of $N$ points. For instance, for the case of $N=3$ (triangle) each angle is taken into account twice by the definition (1), the combinatorial factor $N(N-1)(N-2)=6$ and thus normalizes $\lambda_{3}$. By way of example, $\lambda_{3}=-0.5$ in the case of the symmetrical configuration of three points in a plane (the equilateral triangle). It is important to note that $\lambda_{N}=1$ independently of the number of points under consideration if they lie exactly along the straight line. The degree of alignment $P_{N}$ is understood as the fraction of the events for which $\lambda_{N}>0.8$ book among the total number of events in which the number of energy centres not less than $N$.
The needed azimuthal angles are calculated over the positions ${\bf r}_{i}$ of the particles (spots) in the $(xy)$-plane in the film:
${\bf r}_{i}~{}=~{}\frac{{\bf v}_{ri}}{v_{zi}}~{}h~{}=\frac{{\bf p}_{Ti}}{\sqrt{p^{2}_{Ti}+m^{2}_{i}}~{}\sinh(\eta_{0}+\eta_{i})}~{}h~{},$ (2)
where $v_{zi}$ and ${\bf v}_{ri}$ are the longitudinal and radial components of particle velocity respectively. $h$ is the height of several hundred meters to several kilometers in the atmosphere above chamber where the primary interaction takes place.
For convenience we parametrize the 4-momentum of each particle $i$ under consideration with its transverse momentum $p_{Ti}$ (with respect to the collision axis $z$), pseudorapidity $\eta_{i}$ and azimuthal angle $\phi_{i}$ in the center-of-mass system:
$[\sqrt{p^{2}_{Ti}+m^{2}_{i}}~{}\cosh\eta_{i},~{}~{}~{}p_{Ti}\cos\phi_{i},~{}~{}~{}p_{Ti}\sin\phi_{i},$
$\sqrt{p^{2}_{Ti}+m^{2}_{i}}~{}\sinh\eta_{i}].$ (3)
At the considered interacting energies particles are ultrarelativistic, and, as is known, in this case the values of rapidity and pseudorapidity are practically the same. For this reason, in what follows, under rapidity we understand pseudorapidity. The transformation from the center-of-mass system to the laboratory frame amounts to the rapidity shift: $\zeta_{i}=\eta_{0}+\eta_{i}$, where $\eta_{0}$, $\zeta_{i}$ are the rapidities of the center-of-mass system and the particle $i$ respectively in the laboratory reference frame.
Because the size of the observation region is about several centimeters, these distances $r_{i}$ must satisfy the following relations:
$r_{\rm min}~{}<~{}r_{i},$ (4)
$r_{i}~{}<~{}r_{\rm max}.$ (5)
The condition (4) speaks about noncoincidence of the points in the film with the center formed by the particles emitted along the collision axis (predominantly region of incident-hadron fragmentation). The distance between particles is determined by the expression:
$d_{ij}~{}=~{}\sqrt{r^{2}_{i}~{}+~{}r^{2}_{j}~{}-~{}2r_{i}r_{j}\cos(\phi_{i}~{}-~{}\phi_{j})}$ (6)
and must satisfy the condition:
$d_{ij}~{}>~{}r_{\rm min}~{},$ (7)
which means the distinguishability of spots $i$ and $j$ from each other. Otherwise the particles are combined into a cluster until there remain only particles and/or particle clusters with mutual distances larger than $r_{\rm min}$. The coordinates of the new cluster formed by two particles are determined in just the same way as the center-of-mass coordinates of two bodies in classical mechanics:
${\bf r}_{ij}=({\bf r}_{i}E_{i}+{\bf r}_{j}E_{j})/(E_{i}+E_{j}).$ (8)
Among clusters that satisfy the restrictions (4), (5) and (7) one selects the $2,...,7$ clusters/particles which are most energetic. After that one calculates the alignment $\lambda_{N}$ using the definition above and taking into account the central cluster, i.e. $N-1=2,...,7.$
3 HYDJET++ model
An absence of any collider experiment evidences for exotic types of interactions encouraged us to investigate more intently the influence of the selection procedure of most energetic particles itself and the threshold on the energy deposition in appearing of alignment phenomenon. These studies Lokhtin:2023tze was based on the geometrical and kinematical considerations being not influenced by the specific dynamics. Now we would like spreading this hopeful investigation to some realistic model of hadron interactions. For this purpose we use the popular and known HYDJET++ model which successfully describes the large number of physical observables measured in heavy ion collisions during RHIC and LHC operation. The details on this model may be found in the HYDJET++ manual HYDJET_manual .
The event generator includes two independent components: the soft, hydro-type state and the hard state resulting from in-medium multi-parton fragmentation. The soft component is the “thermal” hadronic state generated on the chemical and thermal freeze-out hypersurfaces prescribed by the parametrization of relativistic hydrodynamics with preset freeze-out conditions (the adapted event generator FAST MC Amelin:2006qe ; Amelin:2007ic ). Particle multiplicities are calculated within the effective thermal volume using a statistical model approach. The effective volume absorbs the collective velocity profile and the form of hypersurface and cancels out in all particle number ratios. Therefore, the particle number ratios do not depend on the freeze-out details as long as the local thermodynamic parameters are independent of spatial coordinates. The concept of the effective volume is applied to calculate the hadronic composition at both chemical and thermal freeze-outs. The number of particles in an event is calculated according to Poisson distribution around its mean value, which is supposed to be proportional to the number of participating nucleons for a given impact parameter in a A+A collision. To simulate the elliptic and triangular flow effects, the hydro-inspired parametrization for the momentum and spatial anisotropy of soft hadron emission source is implemented HYDJET_manual ; Wiedemann:1997cr .
In the hard sector the model propagates the hard partons through the expanding quark gluon plasma and takes into account both collisional loss and gluon radiation due to parton rescattering. It is based on the PYQUEN partonic energy loss model Lokhtin:2005px . The number of jets is generated according to the binomial distribution. Their mean number in an A+A event is calculated as the product of the number of binary nucleon-nucleon (NN) sub-collisions at a given impact parameter and the integral cross section of the hard process in NN collisions with the minimum transverse momentum transfer $p_{T}^{\rm min}$. The latter is the input parameter of the model. In the HYDJET++ framework, partons produced in (semi)hard processes with the momentum transfer lower than $p_{T}^{\rm min}$ are considered as being “thermalized”, so their hadronization products are included in the soft component of the event “automatically”.
It should be noted that applying this model we have the natural threshold on collision energy in appearing of alignment phenomenon since the conception of quark gluon plasma is valid after some minimal colliding energy.
4 Simulation of alignment in HYDJET++ model
As it was mention above the model calculations such as transverse momentum spectra, pseudorapidity and centrality dependence of inclusive charged particle multiplicity and $\pi^{\pm}\pi^{\pm}$ correlation radii in central Pb+Pb collisions Lokhtin:2012re , centrality and momentum dependence of second and higher-order harmonic coefficients Bravina:2013xla , flow fluctuations Bravina:2015sda , jet quenching effects jet1 ; jet2 , and angular dihadron correlations Eyyubova:2014dha are in fair agreement with the experimental data. Therefore, to be specific we consider Pb+Pb collisions in the centrality class $c=0-5\%$ at center-of-mass energy 5.02 TeV per nucleon pair with free parameters of the model tuned and fixed early. Following the consideration in the previous section 2 we set the parameters $r_{\rm min}=r_{\rm res}=1$mm, $r_{\rm max}=15$mm, $h=1$km, which are close to the conditions of emulsion experiments, with the additional restriction on the energy threshold of particle registration in the emulsion $E_{i}>E^{\rm thr}=4$TeV. Our calculations are practically insensitive to this threshold in the wide interval of its variation. The results for the degree of alignment $P_{N}$ are the following:
$P_{3}\approx 0.2,~{}P_{4}\approx 0.04,~{}P_{5}\approx 0.008~{}\rm at~{}\lambda_{N}>0.8.$ (9)
and are close to that obtained early Lokhtin:2023tze ; LokhtinYF24 for chaotically located spots in the X-ray film. It is surprising at first glance only. In fact the azimuthal angle distribution of particles in most central collisions is practically isotropic in the HYDJET++ generator as the same as the angular distribution of points in the procedure applied in Lokhtin:2023tze ; LokhtinYF24 and the existing angular model correlations (for instance, with the reaction plane) are not enough to reproduce the correlations needed for appearing of alignment phenomenon without the special selection procedure.
Figure 1: The degree of alignment $P_{3}$ for the three particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
Figure 2: The degree of alignment $P_{4}$ for the four particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
Figure 3: The degree of alignment $P_{5}$ for the five particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
Here it is worth noting that the majority of soft particles are generated independently in the statistical model approach. In this approach the total momentum-energy together with the particle number vary (fluctuate) from an event to an event. Only their mean values are meaningful that is the transverse momentum of all particles is equal to zero not every event, but in average. To decide this long-standing conceptual problem, the possible influence of transverse momentum conservation in every event we take into account in the form of missing transverse momentum:
$|{\bf p}_{T_{1}}+{\bf p}_{T_{2}}+...+{\bf p}_{T_{(N-1)}}|<\Delta,$ (10)
where ${\bf p}_{T_{i}}$ is the transverse momentum of $i$ particle. The smaller the value $\Delta$ the better the transverse momentum of the selected highest-energy particles should be balanced and vice versa – the larger $\Delta$ the greater the disbalance of the transverse momentum is permissible for the selected particles.
For the two selected most energetic particles the value of their total transverse momentum $\Delta_{2}$ is equal to
$\Delta_{2}=\sqrt{p^{2}_{T1}+p^{2}_{T2}+2p_{T1}p_{T2}\cos{(\phi_{1}-\phi_{2})}}.$ (11)
In the case of the collinear ($\phi_{1}-\phi_{2}\simeq 0$) and back-to-back ($|\phi_{1}-\phi_{2}|\simeq\pi$) configurations one has the “line up events” with the large value of the parameter $\lambda_{3}$. The share of such configurations among all possible ones is just $\sim 20\%$ and can be directly estimated if the azimuthal angular distribution of particles is known.
With the decreasing of $\Delta$ the condition (10)( $\Delta_{2}<\Delta$) strengthens the fraction of back-to-back configurations, since they have the minimal value of $\Delta_{2}\simeq|p_{T1}-p_{T2}|$ among its possible values. The condition (4) about noncoincidence of the points in the film with the center means also that there is some minimal transverse momentum $p_{T{\rm min}}$ of selected highest-energy particles in according with the kinematical relation (2). This minimal transverse momentum depends on the size ($r_{\rm min}$) of the region observation, the height ($h$) of primary interaction and the collision energy (over $\eta_{0}$). In fact it characterises the scale of missing transverse momentum and the some measure of transverse momentum resolution. Thus the restriction on the transverse momentum from below means that at $\Delta<\sqrt{2}p_{T{\rm min}}$ survives back-to-back configurations only and we have practically the $100\%$-degree of alignment. This peculiarity in the behavior of $P_{3}$ as a function of transverse momentum disbalance $\Delta$ is shown in Fig.1. In according with the kinematical relation (2) the region of high degree of alignment becomes wider with the height decreasing, that is also seen in Figs.1a–b, because in this case the minimal transverse momentum becomes larger too. However this widening is not so large since the inverse proportionality between the transverse momentum and the height in the numerator of (2) is partially compensated by the energy decreasing in the denominator.
The results of our modelling with restriction (10) are shown in Fig.2 and Fig.3 for the four and five of selected highest-energy particles respectively. The degree of alignment increases noticeably by a factor of the order of 2 for the $P_{4}$ and a factor of order 5 for the $P_{5}$ in comparison with its values (9) without restriction (10). The effect of the odd number of considered particles takes place also and which is observed in our previous investigation Lokhtin:2023tze ; LokhtinYF24 at the odd number of the points randomly generated. The reason is similar: the disbalance $\Delta$ for the three selected particles is of the order of $p_{T{\rm min}}$ or larger, if their location is close to the same line passing through an origin (center). The small value of $\Delta$ does not allow such configurations to be generated and the degree of alignment is close to zero.
To disentangle the influence of different mechanics of particle production we considered at first only the soft part of the HYDJET++ model (“without jets” in Figs.13), that is the adapted version of event generator FASTMC and which yields the bulk of particles with the transverse momentum up to $\sim 5$ GeV. In the interval of the transverse momentum larger than $\sim 5$ GeV the influence of hard part of the considered model can be noticeable.
Figures1c–3c show the compare of the results with included jet mechanism (“with jets”) and with only soft particles. The form of curves is not practically changed, but significant increasing of alignment degree takes place for the four and five particles analyzed. Moreover the effect of odd particle number is less pronounce, since the small disbalance is achievable for the back-to back configurations with the large enough value of $\lambda$, because the transverse momentum of one jet particle can be easily compensated by the total transverse momentum of two soft particles having relatively less characteristic momentum.
To investigate the possible influence of elliptic and triangular flow effects (included in the event generating mode) on the alignment degree $P_{N}$ we consider also so-called “minimum bias”, in fact centrality class collisions $c=0-75\%$, Figs.1b–3b. It is very difficult to assess the effect on alignment from an anisotropic flow with the jet mechanism turned on, since the contribution of jets to the full multiplicity depends significantly on centrality and this introduces additional uncertainty into the simulation result, so it is reasonable to compare the results for $P_{N}$ only for soft particles (without jets). The results for $P_{3}$ and $P_{5}$ remain practically unchanged, but for $P_{4}$ we observe less pronounce effect of the odd number of considered particles as it shown in Fig.2b. The explanation is close to the case with jets mechanism: the small disbalance $\Delta$ is achievable for the back-to back configurations with the large enough value of $\lambda$, because the transverse momentum of one particle strengthen by anisotropic flow can be easily compensated by the total transverse momentum of two soft particles having relatively less characteristic momentum.
Thus our modeling demonstrates clearly that the conservation law of the transverse momentum in the form of the disbalance of the total transverse momentum $\Delta$ of highest-energy particles allows one to select automatically more line up configurations. The application of this procedure to the HYDJET++ model can be partially justified by the fact that the majority of soft particles are generated independently as it was mention above and the transverse momentum conserves only in average, but in any rapidity interval. Conceptually similar modification of the statistical approach has been recently done Chernyshov_2023 to take into account the charge conservation in event-by-event to describe successfully the LHC data on the charge balance function. The high degree of alignment $P_{3}$ is achievable at the reasonable values (up to $\sim 3$ GeV) of transverse momentum disbalance of selected most energetic particles.
It worth noting that the degree of alignment $P_{N}$ is not enough large to match the central values of experimental measurements for $N\geq 4$. For comparison the Pamir measurements are $P_{4}^{\rm exp}=0.67\pm 0.33$ and $P_{5}^{\rm exp}=0.33\pm 0.23$ man , while our “best” values are $P_{4}\approx 0.07$ and $P_{5}\approx 0.04$. Nevertheless these values are in the limits of the two standard deviations from the experimentally measured values as a minimum.
5 Conclusions
We have analyzed the influence of the selection procedure of most energetic particles itself in appearing of alignment phenomenon in the framework of HYDJET++ model. Event-by-event the transverse momentum conversation has been taken into account in the form of missing transverse momentum since the majority of particles are independently generated in the statistical model approach. It has been shown that the high degree of alignment appears at the reasonable values (up to $\sim 3$ GeV) of transverse momentum disbalance of selected most energetic particles. Applying this model we have the natural threshold on collision energy in appearing of alignment phenomenon since the conception of quark gluon plasma is valid after some minimal colliding energy. However one should note that the degree of alignment $P_{N}$ in our modeling is not enough large to match the central values of experimental measurements for $N\geq 4$ living some room for other explanations and further investigations.
Acknowledgements.
It is pleasure to thank A.S. Chernyshov, A.I. Demianov, S.N. Nedelko for fruitful discussions. This work is supported by the Russian Science Foundation, grant 24-22-00011.