### v3

# Unraveling alignment pattern in high-energy particles via transverse momentum disbalance analysis

Abstract
The hypothesis of the relation between the strong azimuthal correlations, known as the alignment, in the families of hadrons and photons observed in cosmic ray experiments and the selection procedure of the highest-energy particles together with the transverse momentum conservation are tested in the framework of the HYDJET++ model. Results show that the high degree of alignment can appear in nucleus-nucleus interactions at the reasonable values of transverse momentum disbalance of most energetic detected particles.
Properties of Hadrons; Specific QCD Phenomenology; Relativistic Heavy Ion Physics; Particle Correlations and Fluctuations; Quark-Gluon Plasma; Cosmic Rays
I Introduction
In our previous study [1, 2] in the framework of geometrical approach it has been shown that the high degree of alignment can appear partly due to the selection procedure of the highest-energy particles itself and the threshold on the energy deposition together with the transverse momentum conservation. This hopeful result encourages us to investigate more intently this hypothesis in the framework of some realistic model of hadron interactions. It is about the observation of strong angular correlations in the families of hadrons and photons that are products of interaction between cosmic rays and target nuclei [3, 4, 5, 6, 7]. This phenomenon can potentially indicate the coplanar structure of the observed events and was named alignment [6]. Among other angular characteristics, alignment most exactly demonstrates the position of the most energetic registered particles (or their clusters) relative to a straight line in the plane of the emulsion film. In addition to that the strong angular correlations of particles were also observed in stratospheric [8, 9] X-ray-emulsion chamber experiments. So far there is no commonly accepted explanation of this intriguing phenomenon in spite of numerous attempts to find it (see, for instance, [7, 10, 11, 12, 13, 14, 15] and references therein).
The long-range azimuthal correlations, also known as the ridge effect, made in heavy ion collisions at the RHIC [16] and in $pp$-interactions with high multiplicity [17] at the LHC, stimulated the search of any manifestations [18, 19, 20] of alignment phenomenon at the LHC conditions. Presumably, the alignment should already appear at the LHC since the total collision energy of the beams in these experiments exceeds the energy threshold $\sqrt{s_{\rm eff}}\geq 4$ TeV for detected particles with alignment in experiments with cosmic rays. However, the determination of any correspondence between these striking phenomena was unsuccessful because they were observed in the quite different intervals of the rapidity practically without overlapping and in the different reference frames. Moreover the subsequent studies allowed finding the explanation of the ridge effect within established theoretical approaches: For instance, as a simple interplay between the elliptic and triangular flows [21]. This fact basically agrees with a rather widespread point of view (questioned in [12, 22, 23]) that the alignment phenomenon is not more than a tail in a distribution caused fluctuations. In this paper we present a development and continuation of the ideas described in our recent studies [1, 2], but our approach is now based on the well-known and successfully tested model of heavy-ion collisions HYDJET++  [24]. In current paper, we consider the modification of the statistical approach in the model which conceptually is similar to that proposed recently in [25] to take into account the event-by-event conservation of the electric net-charge of produced particles. This allowed us to reproduce the LHC data on charge balance functions measured in Pb+Pb collisions. The paper is organized as follows: in Sect. 2 we remind the main definitions and kinematical relations in the problem under consideration. Basic characteristics of the applied HYDJET++ model are sketched in Sect. 3. The simulation results and their description are presented in Section 4. Section 5 is devoted to a summary, conclusions and discussion.
II Basic definitions and kinematical relations
As mentioned above, the alignment was observed in experiments with high-energy cosmic rays, so all kinematic, geometric, angular relations of the detected particles in the plane of the emulsion film are related to the laboratory coordinate system. To characterize the azimuthal angular correlations the Pamir Collaboration defines the alignment of spots  [6] in term of the variable
$\lambda_{N}~{}=~{}\frac{\sum^{N}_{i\neq j\neq k}\cos(2\varphi_{ijk})}{N(N-1)(N-2)},$ (1)
where $\varphi_{ijk}$ is the angle between two straight lines connecting the $i$spot with the the $j$th and $k$th spots. $N$ is the number of spots used in an analysis. The alignment, $\lambda_{N}$, quantitatively describes a deviation of points from the straight line and qualitatively characterized the asymmetry better than the possible other parameters like the eccentricity or the thrust. For instance, $\lambda_{4}$ will be equal to 1 if all four points lie strictly on the same straight line, but it will be considerably less than 1 if these points form four vertices of a long rectangle. The combinatorial factor $(N(N-1)(N-2))$ normalizes the parameter $\lambda_{N}$, because it takes into account the number of variants of choosing three of $N$ points. In the case, example, of $N=3$, according the definition  (1) each angle of triangle is taken into account twice, the combinatorial factor $N(N-1)(N-2)=6$ and thus normalizes $\lambda_{3}$. By way of example, $\lambda_{3}=-0.5$ in the case of the symmetrical configuration of three points in a plane (the equilateral triangle). It is important to note that $\lambda_{N}=1$ independently of the number of points under consideration if they lie exactly along the straight line. The ratio of the number of events for which $\lambda_{N}>0.8$ to the total number of events (in which the number of energy centres not less than $N$) is called the degree of alignment $P_{N}$.
The needed azimuthal angles are calculated over the positions ${\bf r}_{i}$ of the particles (spots) in the $(xy)$-plane in the film:
${\bf r}_{i}~{}=~{}\frac{{\bf v}_{ri}}{v_{zi}}~{}h~{},$ (2)
where $v_{zi}$ and ${\bf v}_{ri}$ are the longitudinal and radial components of particle velocity respectively. $h$ is the altitude of several hundred meters to several kilometers in the Earth atmosphere above emulsion chamber where the primary interaction with cosmic rays takes place. The size of particle observation is about several centimeters and related to the characteristics of the X-ray film used in the Pamir experiment. The gap for $r_{i}$ must satisfy the following relation
$r_{\rm min}~{}<~{}r_{i}~{}<~{}r_{\rm max}.$ (3)
The condition (3) speaks about noncoincidence of the points in the film with the center $r_{\rm min}$, that designed by the different particles moving along the collision axis (mainly region of incident-hadron dissolution), and location of spots inside observation film sector. The distance between particles is determined by the expression:
$d_{ij}~{}=~{}\sqrt{r^{2}_{i}~{}+~{}r^{2}_{j}~{}-~{}2r_{i}r_{j}\cos(\phi_{i}~{}-~{}\phi_{j})}$ (4)
$\phi_{i}$ are the the azimuthal angle and must satisfy the condition:
$d_{ij}~{}>~{}r_{\rm min}~{},$ (5)
which means the distinguishability of spots $i$ and $j$ from each other. Otherwise the particles are combined into a cluster until there remain only particles and/or particle clusters with mutual distances larger than $r_{\rm min}$. The new cluster that is created by two particles has the position at the emulsion film
${\bf r}_{ij}=({\bf r}_{i}E_{i}+{\bf r}_{j}E_{j})/(E_{i}+E_{j})$ (6)
and managed as the center-of-mass coordinates of two bodies in classical mechanics, where $E_{i},~{}E_{j}$ is energy of $i$ and $j$ particles.
The alignment is calculated by the expression (1), where except the central bunch in between clusters that satisfy the restrictions (3), (5) one selects the $2,...,7$ clusters/particles which are most energetic, i.e. $N-1=2,...,7.$
III HYDJET++ model
Since no manifestation of the alignment presence have been obtained for the entire time of collider experiments, and also in the available data from cosmic ray experiments there is a large measurement error, in our recent papers  [1, 2] we tried to understand how the selection procedure of the most energetic particles together with the threshold of its energy influence at the pattern of alignment. These studies based on the geometrical and kinematical considerations being not influenced by the specific dynamics. Now we would like spreading this hopeful investigation to some realistic model of hadron interactions. For this purpose we use the popular and known HYDJET++ model which successfully describes the large number of physical observables measured in heavy ion collisions during RHIC and LHC operation. The details on this model may be found in the HYDJET++ manual [24].
The event generator includes two independent components: the soft, hydro-type state and the hard state resulting from in-medium multi-parton fragmentation. The soft component is the “thermal” hadronic state generated on the chemical and thermal freeze-out hypersurfaces prescribed by the parametrization of relativistic hydrodynamics with preset freeze-out conditions (the adapted event generator FAST MC [26, 27]). Particle multiplicities are calculated within the effective thermal volume using a statistical model approach. The effective volume absorbs the collective velocity profile and the form of hypersurface and cancels out in all particle number ratios. Therefore, the particle number ratios do not depend on the freeze-out details as long as the local thermodynamic parameters are independent of spatial coordinates. The concept of the effective volume is applied to calculate the hadronic composition at both chemical and thermal freeze-outs. The number of particles in an event is calculated according to Poisson distribution around its mean value, which is supposed to be proportional to the number of participating nucleons for a given impact parameter in a A+A collision. To simulate the elliptic and triangular flow effects, the hydro-inspired parametrization for the momentum and spatial anisotropy of soft hadron emission source is implemented [24, 28].
In the hard sector the model propagates the hard partons through the expanding quark gluon plasma and takes into account both collisional loss and gluon radiation due to parton rescattering. It is based on the PYQUEN partonic energy loss model [29]. The number of jets is generated according to the binomial distribution. Their mean number in an A+A event is calculated as the product of the number of binary nucleon-nucleon (NN) sub-collisions at a given impact parameter and the integral cross section of the hard process in NN collisions with the minimum transverse momentum transfer $p_{T}^{\rm min}$. The latter is the input parameter of the model. In the HYDJET++ framework, partons produced in (semi)hard processes with the momentum transfer lower than $p_{T}^{\rm min}$ are considered as being “thermalized”, so their hadronization products are included in the soft component of the event “automatically”.
It should be noted that applying this model we have the “natural” threshold on collision energy in appearing of alignment phenomenon since the conception of quark gluon plasma is valid after some minimal colliding energy.
IV Simulation of alignment in HYDJET++ model
As noted in the previous section the model calculations such as transverse momentum spectra, pseudorapidity and centrality dependence of inclusive charged particle multiplicity and $\pi^{\pm}\pi^{\pm}$ correlation radii in central Pb+Pb collisions [30], centrality and momentum dependence of second and higher-order harmonic coefficients [31], flow fluctuations [32], jet quenching effects [33, 34], and angular dihadron correlations [21] are in fair agreement with the experimental data. Therefore, to be specific we consider Pb+Pb collisions in the centrality class $c=0-5\%$ at center-of-mass energy 5.02 TeV per nucleon pair with free parameters of the model tuned and fixed early. Of course, Pb+Pb system is much larger than one created in interactions of cosmic nuclei in the atmosphere, like Fe+O reactions, but it is available for studying in current collider (LHC) experiments. We suppose that alignment phenomenon can be assisted nucleus-nucleus interactions at high energies, including some collective effects together with local transverse momentum conservation and so may studied in collider as well as in cosmic ray experiments.
Since the alignment was observed in the emulsion film, to that the laboratory frame is bound, but we are modeling the collisions of relativistic heavy nuclei in the center-of-mass system, it is convenient to write the 4-momentum of each particle $i$ under study as follows
$[\sqrt{p^{2}_{Ti}+m^{2}_{i}}~{}\cosh\eta_{i},~{}~{}~{}p_{Ti}\cos\phi_{i},~{}~{}~{}p_{Ti}\sin\phi_{i},$
$\sqrt{p^{2}_{Ti}+m^{2}_{i}}~{}\sinh\eta_{i}],$ (7)
where $p_{Ti}=\sqrt{p_{x_{i}}^{2}+p_{y_{i}}^{2}}$ is the transverse momentum (with respect to the collision axis $z$), $\eta_{i}$, $\phi_{i}$ are the pseudorapidity and the azimuthal angle in the center-of-mass system. At the considered interacting energies particles are ultrarelativistic, and, as is known, in this case the values of rapidity and pseudorapidity are practically the same. For this reason, in what follows, under rapidity we understand pseudorapidity. An important element of our simulation is the transformation from the center-of-mass system to the laboratory frame that realized by the rapidity shift: $\zeta_{i}=\eta_{0}+\eta_{i}$, where $\eta_{0}$, $\zeta_{i}$ are the rapidities of the center-of-mass system and the particle $i$ respectively in the laboratory reference frame. The positions $r_{i}$ of particles with mass $m_{i}$ reads as
${\bf r}_{i}=\frac{{\bf p}_{T_{i}}}{\sqrt{p_{T_{i}}^{2}+m_{i}^{2}}\sinh(\eta_{0}+\eta_{i})}~{}h.$ (8)
Following the consideration in the previous section 2 we set the parameters $r_{\rm min}=r_{\rm res}=1$mm, $r_{\rm max}=15$mm, $h=1$km, which are close to the conditions of emulsion experiments, with the additional restriction on the energy threshold of particle registration in the emulsion $E_{i}>E^{\rm thr}=4$TeV. Our calculations are practically insensitive to this threshold in the wide interval of its variation. The results of simulation for the degree of alignment $P_{N}$ in the framework of HYDJET++ model are the following:
$P_{3}\approx 0.2,~{}P_{4}\approx 0.04,~{}P_{5}\approx 0.008~{}\rm at~{}\lambda_{N}>0.8$ (9)
and are close to that obtained early [1, 2] for chaotically located spots in the X-ray film. It is surprising at first glance only. In fact the azimuthal angle distribution of particles in most central collisions is practically isotropic in the HYDJET++ generator as the same as the angular distribution of points in the procedure applied in [1, 2] and the existing angular model correlations (for instance, with the reaction plane) are not enough to reproduce the correlations needed for appearing of alignment phenomenon without the special selection procedure.
Here it is worth noting that the majority of soft particles are generated independently in the statistical model approach. In this approach the total momentum-energy together with the particle number vary (fluctuate) from an event to an event. Only their mean values are meaningful that is the transverse momentum of all particles is equal to zero not every event, but in average. To decide this long-standing conceptual problem, the possible influence of transverse momentum conservation in every event we take into account in the form of missing transverse momentum:
$|{\bf p}_{T_{1}}+{\bf p}_{T_{2}}+...+{\bf p}_{T_{(N-1)}}|<\Delta,$ (10)
where ${\bf p}_{T_{i}}$ is the transverse momentum of $i$ particle. The smaller the value of $\Delta$, the better the transverse momentum of the selected highest-energy particles should be balanced and vice versa – the larger $\Delta$, the greater the disbalance of the transverse momentum is permissible for the selected particles.
Figure 1: The degree of alignment $P_{3}$ for the three particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
For the two selected most energetic particles (or clusters) and one particle in the origin ($N=3$) the value of its total transverse momentum $\Delta_{2}$ is equal to
$\Delta_{2}=\sqrt{p^{2}_{T1}+p^{2}_{T2}+2p_{T1}p_{T2}\cos{(\phi_{1}-\phi_{2})}}.$ (11)
In the case of the collinear ($\phi_{1}-\phi_{2}\simeq 0$) and back-to-back ($|\phi_{1}-\phi_{2}|\simeq\pi$) configurations one has the “line up events” with the large value of the parameter $\lambda_{3}$. The share of such configurations among all possible ones is just $\sim 20\%$ and can be directly estimated if the azimuthal angular distribution of particles is known.
Figure 2: The degree of alignment $P_{4}$ for the four particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
Figure 3: The degree of alignment $P_{5}$ for the five particles as a function of the disbalance $\Delta$. (a) – the results for the only soft particles (“without jets”) at the different values of the height $h$ in the centrality class collisions $c=0-5\%$; (b) – the same the as (a), but in the centrality class collisions $c=0-75\%$; (c) – comparing results for the soft particles and the case with included jet mechanism (“with jets”) at the height $h=1$km.
With the decreasing of $\Delta$ the condition (10)( $\Delta_{2}<\Delta$) strengthens the fraction of back-to-back configurations, since they have the minimal value of $\Delta_{2}\simeq|p_{T1}-p_{T2}|$ among its possible values. The condition (3) about noncoincidence of the points in the film with the center means also that there is some minimal transverse momentum $p_{T{\rm min}}$ of selected highest-energy particles in according with the kinematical relation (2). This minimal transverse momentum depends on the size ($r_{\rm min}$) of the region observation, the height ($h$) of primary interaction and the collision energy (over $\eta_{0}$). In fact it characterises the scale of missing transverse momentum and the some measure of transverse momentum resolution. Thus the restriction on the transverse momentum from below means that at $\Delta<\sqrt{2}p_{T{\rm min}}$ survives back-to-back configurations only and we have practically the $100\%$-degree of alignment. This peculiarity in the behavior of $P_{3}$ as a function of transverse momentum disbalance $\Delta$ is shown in figure 1. In according with the kinematical relation (2) the region of high degree of alignment becomes wider with the height decreasing, that is also seen in the figures 1a–b, because in this case the minimal transverse momentum becomes larger too. However this widening is not so large since the inverse proportionality between the transverse momentum and the height in the numerator of (2) is partially compensated by the energy decreasing in the denominator.
The results of our modelling with restriction (10) are shown in figure 2 and figure 3 for the four and five of selected highest-energy particles respectively. The degree of alignment increases noticeably by a factor of the order of 2 for the $P_{4}$ and a factor of order 5 for the $P_{5}$ in comparison with its values (9) without rule (10). The effect of the odd quantity particles takes place also for $P_{4}$ and which is observed in our previous studies [1, 2] at an odd number of the points randomly generated. The reason is similar: the disbalance $\Delta$ for the three selected particles is of the order of $p_{T{\rm min}}$ or larger, if their location is close to the same line passing through an origin (center). The small value of $\Delta$ does not allow such configurations to be generated and the degree of alignment $P_{4}$ is close to zero.
To disentangle the influence of different mechanics of particle production we considered at first only the soft part of the HYDJET++ model (“without jets” in figures 13), that is the adapted version of event generator FASTMC and which yields the bulk of particles with the transverse momentum up to $\sim 5$ GeV. In the interval of the transverse momentum larger than $\sim 5$ GeV the influence of hard part of the considered model can be noticeable.
Figures1c–3c show the compare of the results with included jet mechanism (“with jets”) and with only soft particles. The form of curves is not practically changed, but significant increasing of alignment degree takes place for the four and five particles analyzed. Moreover the effect of odd particle number is less pronounce for $P_{4}$, since the small disbalance $\Delta$ is achievable for the back-to back configurations with the large enough value of $\lambda$, because the transverse momentum of one jet particle can be easily compensated by the total transverse momentum of two soft particles having relatively less characteristic momentum.
To investigate the possible influence of elliptic and triangular flow effects (included in the event generating mode) on the alignment degree $P_{N}$ we consider also so-called “minimum bias”, in fact centrality class collisions $c=0-75\%$, figures 1b–3b. In addition, at $c=0-75\%$ the modeling parameters are as close as possible to the conditions of experiments with cosmic rays. It is challenging to evaluate the impact of the anisotropic flow on alignment with the jet mechanism activated, as the contribution of jets to the overall multiplicity significantly depends on centrality, which introduces additional uncertainty in the simulation results. Therefore, it would be reasonable to compare the results of $P_{N}$ for soft particles only, i.e. without jets. The results for $P_{3}$ and $P_{5}$ remain practically unchanged, but for $P_{4}$ we observe less pronounce effect of the odd quantity of considered particles as it shown in figure 2b. We associate this with the manifestation of the transverse momentum equality, enhanced by an anisotropic flow, of one particle with the total transverse momentum of two “softer” particles, each of which has a relatively smaller characteristic momentum. Therefore, high alignment values can only be realized with relatively low levels of missing momentum $\Delta$.
V Discussion and Conclusions
Presumably, the alignment can manifest itself in both high energy collisions of relatively light nuclei (in cosmic rays experiments) and heavy ions (in collider experiments) together with collective effects, such as quark-gluon plasma formation and the local conservation of the transverse momentum of detected particles and their clusters. We have analyzed the influence of the selection procedure of most energetic particles itself in appearing of alignment phenomenon in the framework of HYDJET++ model. Event-by-event the transverse momentum conservation has been taken into account in the form of missing transverse momentum since the majority of particles are independently generated in the statistical model approach. The results on the figures 13 show that the high degree of alignment appears at the reasonable values (up to $\sim 3$ GeV) of transverse momentum disbalance of selected most energetic particles. In addition, the HYDJET++ model assumes the statistical particle production due to the hadronization of the quark-gluon plasma that automatically implies a collision energy threshold. This threshold is also characteristic of alignment observation.
Figure 4: Comparison of the simulation results of the alignment degree $P_{N}$ for three, four, five particles with the data of the Pamir collaboration. (a) – linear scale, (b) – logarithmic scale. “HJ++ with $\Delta$” – the modeling with taking account the local transverse momentum conservation with disbalance $\Delta$ (10); “HJ++” – the modeling without ${\bf p}_{T}$ conservation in every event. The values of the disbalance $\Delta$, which correspond to the our “best” alignment degree, are in the range of $0\div 1$ GeV.
A comparison of the simulation results with the Pamir collaboration data is shown in figure 4. It should be noted that since the Pamir data were obtained from experiments with cosmic rays, while our results are related to heavy ion collisions, rather an qualitative comparison can be made. Figure 4 shows that the alignment degree of three particles ($N=3$) can be fully described within the HYDJET++ model taking into account the local transverse momentum conservation with the disbalance $\Delta$, whereas for four ($N=4$) and five ($N=5$) particles we do not reproduce the experimental data [11] even taking into account measurement error:
$P_{3}^{\rm exp}=0.83\pm 0.27;~{}P_{4}^{\rm exp}=0.67\pm 0.33;~{}P_{5}^{\rm exp}=0.33\pm 0.23,$ (12)
while our “best” values are
$P_{3}=1;~{}P_{4}\approx 0.07;~{}P_{5}\approx 0.04,$ (13)
which achieved at disbalance in the interval
$\Delta=0\div 1~{}\text{GeV},$ (14)
corresponding to the transverse momentum conservation with high enough accuracy. Nevertheless the alignment degree $P_{4}$ and $P_{5}$ are in the limits of the two standard deviations from the experimentally measured values as a minimum, in addition, the values $P_{4}$ and $P_{5}$ increase by a factor of about 2 and 5, respectively, compared to result (9) due to ${\bf p}_{T}$ conservation in every event.
Thus our modeling demonstrates clearly that the conservation law of the transverse momentum via minimization of the disbalance $\Delta$ of the total transverse momentum of highest-energy particles allows one to select automatically more line up configurations. The application of this procedure to the HYDJET++ model can be partially justified by the fact that the majority of soft particles are generated independently as it was mention above and the transverse momentum conserves only in average, but in any rapidity interval. However one should note that the degree of alignment $P_{N}$ in our modeling is not enough large to match the central values of experimental measurements for $N\geq 4$ living some room for other explanations and further investigations.
Acknowledgements.
It is pleasure to thank A.S. Chernyshov, A.I. Demianov for fruitful discussions and S.N. Nedelko for the our communication . This work is supported by the Russian Science Foundation, grant 24-22-00011.